If 5 times the smaller of two numbers is subtracted from twice the larger the result is 16. If the larger is increased by 3 times the smaller the result is 63. Please help me find the numbers. I put as my let statements: let k=the smaller number;let L=
the larger number

Answer by ankor@dixie-net.com(22740)


write an equation for each statement.
:
If 5 times the smaller of two numbers is subtracted from twice the larger the result is 16.
2L - 5k = 16
If the larger is increased by 3 times the smaller the result is 63.
L + 3k = 63
L = 63 - 3k; use this form for substitution in the 1st equation
:
Replace L in the 1st equation
2(63-3k) - 5k = 16
126 - 6k - 5k = 16
126 - 11k = 16
126 - 16 = 11k
110 = 11k
k = 110/11
k = 10, the smaller number
find k using L = 63 - 3k
L = 63 - 3(10)
L = 63 - 30
L = 33, the larger number
:
You can check the solutions in the 1st statement